OpenStudy (anonymous):

If g(x) = |x-4| and h(x) = x^3; find (h 0 g)(-2) I have no idea what im doing...

hero (hero):

g(x) = |x-4| h(x) = x^3 h o g (x) = h(g(x)) = |x-4|^3 h(g(-2)) = |-2-4|^3 = |-6|^3 = 6^3 = 216

hero (hero):

h(g(x)) means to take g(x) and put it into the input of h(x)

OpenStudy (anonymous):

Thanksss

OpenStudy (anonymous):

g()x=x-4 when x>=4 and g(x) = 4-x when x<4 to find hog(-2) u have to g(x)= 4-x since x=-2<4 h(g(x))=(4-x)^3 at x=-2 we have (4+2)^3=216